how much will $1000 deposited in a single savings account coming in a compound annual interest rate of 6 percent be worth at the end of the following number of years? 3 years
Answer #1


solution FV Future value PV Present value FV=PV*(1+i)An FV=1000* (1.06^ 3) FV Annual Rate of return 1,191.02 number of years